# Deep Automated Mechanism Design for Integrating Ad Auction and Allocation in Feed

### 03 Jan 2024 v1

@@ Abstract.
ccs: Information systems Computational advertisingccs: Information systems Online advertisingccs: Information systems Electronic commerce
@@ 3.1. Setting for Ad Auction and Allocation
$\text{Rev}(b;e)$ Platform expected revenue
$\text{Gmv}(b;e)$ Platform expected GMV
@@ 3.2. Problem Formulation
  • $\mathcal{R}(b;e):\mathcal{B}^{n}\times\mathcal{E}^{m}\to\mathcal{A}$ is the ad allocation rule, which is used to select one winning ad from the $n$ advertisers and insert it at the certain position in the organic items list.
  • $\mathcal{P}(b;e)$ is the payment rule and $p_{i}(b;e)$ is the pay-per-click price of advertiser $i$.
@@
(1) $u_{i}^{\mathcal{M}}(v_{i};b;e)=(v_{i}-p_{i}(b;e))\times q_{\sigma(i)}(\mathcal{R}(b;e)),$
@@
(2) $\text{Rev}^{\mathcal{M}}(b;e)=$ $\sum_{i=1}^{n}p_{i}(b;e))\times q_{\sigma(i)}(\mathcal{R}(b;e)),$
$\text{Gmv}^{\mathcal{M}}(b;e)=$ $\sum_{j=1}^{m}g_{j}(\mathcal{R}(b;e))\times q_{j}(\mathcal{R}(b;e)).$
@@
(3) $u_{i}(v_{i};v_{i},\textbf{b}_{-1};e)\geq u_{i}(v_{i};b_{i},\textbf{b}_{-1};e),\forall b_{i}\in\mathcal{B},$
@@
(5) $\operatorname*{\text{maximize}}_{\mathcal{M}}\quad$ $\mathbb{E}_{\mathbf{a}\in\mathcal{A}}[\text{Rev}(b;e)+\alpha\text{Gmv}(b;e)],$
@@ 5. Experiments
5.1. Experimental Settings
5.1.1. Dataset
Dataset Requests Users Items Avg items
Avito 112,159,463 4,284,824 36,893,299 3.64
Meituan 232,171,262 65,459,719 3,889,579 4.74
@@ 5.2.2. Performance Comparison of mechanisms (Q2)
From the experiment results, we can see that MIAA achieves obtains the highest revenue and GMV, outperforming all baseline mechanisms. Compared with GSP and Fixed Position, MIAA achieves significant improvements, as items with higher eCPM or GMV obtain better positions, and the payment of ads is increased. Compared with GSP and Cross DQN, MIAA improves pRPM. And compared with Score-Weighted VCG, MIAA improves pGPM. Compared with IAS, IMAA, which considers externalities, achieves higher pGMV.

### 11 Apr 2024 v2

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USA.booktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington, DC, USAisbn: 979-8-4007-0431-4/24/07doi: 10.1145/3626772.3657774ccs: Information systems Computational advertisingccs: Information systems Display advertisingccs: Information systems Learning to rank
@@ 3.1. Setting for Ad Auction and Allocation
[…]
@@ 3.2. Problem Formulation
  • $\mathcal{R}(\textbf{b};\textbf{e}):\mathcal{B}^{n}\times\mathcal{E}^{m}\to\mathcal{A}$ is the ad allocation rule, which is used to select one winning ad from the $n$ advertisers and insert it at the certain position in the organic items list.
  • $\mathcal{P}(\textbf{b};\textbf{e})$ is the payment rule and $p_{i}(\textbf{b};\textbf{e})$ is the pay-per-click price of advertiser $i$.
@@
(1) $u_{i}^{\mathcal{M}}(v_{i};\textbf{b};\textbf{e})=(v_{i}-p_{i}(\textbf{b};\textbf{e}))\times q_{\sigma(i)}(\mathcal{R}(\textbf{b};\textbf{e})),$
@@
(2) $\text{Rev}^{\mathcal{M}}(\textbf{b};\textbf{e})=$ $\sum_{i=1}^{n}p_{i}(\textbf{b};\textbf{e}))\times q_{\sigma(i)}(\mathcal{R}(\textbf{b};\textbf{e})),$
$\text{Gmv}^{\mathcal{M}}(\textbf{b};\textbf{e})=$ $\sum_{j=1}^{m}g_{j}(\mathcal{R}(\textbf{b};\textbf{e}))\times q_{j}(\mathcal{R}(\textbf{b};\textbf{e})).$
@@
(3) $u_{i}(v_{i};v_{i},\textbf{b}_{-1};\textbf{e})\geq u_{i}(v_{i};b_{i},\textbf{b}_{-1};\textbf{e}),\forall b_{i}\in\mathcal{B},$
@@
(5) $\operatorname*{\text{maximize}}_{\mathcal{M}}\quad$ $\mathbb{E}_{\mathbf{a}\in\mathcal{A}}[\text{Rev}(\textbf{b};\textbf{e})+\alpha\text{Gmv}(\textbf{b};\textbf{e})],$
@@ 4.4. Training and Online Serving
Dataset Requests Users Items Avg items
Avito 112,159,463 4,284,824 36,893,299 3.64
Meituan 232,171,262 65,459,719 3,889,579 4.74
@@ 5.2.2. Performance Comparison of mechanisms (Q2)
From the experiment results, we can see that MIAA achieves obtains the highest revenue and GMV, outperforming all baseline mechanisms. Compared with GSP and Fixed Position, MIAA achieves significant improvements, as items with higher eCPM or GMV obtain better positions, and the payment of ads is increased. MIAA, considering the incentive effect of organic items on ads, achieves higher pRPM compared with GSP and Cross DQN. Score-Weighted VCG achieves maximum pRPM without considering the loss of GMV. Compared with IAS, MIAA, which considers externalities, achieves higher pRPM and pGMV.